OpenStudy (anonymous):

Can someone help me with Radical Equations? √3x+1=7 Find "x"

OpenStudy (swissgirl):

\( \sqrt{3x+1}=7\) \((\sqrt{3x+1})^2=7^2\) \(3x+1=49\) \(3x=48\) \(x=16\)